OpenStudy (anonymous):

logx – 1 16 = 4 solve for x

OpenStudy (anonymous):

the x-1 is the base

OpenStudy (mertsj):

\[\log_{x-1}16=4 \] \[\log_{x-1}2^4=2^2 \] \[(x-1)^4=2^4\] \[x-1=2\] \[x=3\]